# Small Language Models can Outperform Humans in Short Creative Writing: A Study Comparing SLMs with Humans and LLMs

### v1

@@ Abstract
In this paper, we evaluate the creative fiction writing abilities of a fine-tuned small language model (SLM), BART Large, and compare its performance to humans and two large language models (LLMs): GPT-3.5 and GPT-4o. Our evaluation consists of two experiments: (i) a human evaluation where readers assess the stories generated by the SLM compared to human-written stories, and (ii) a qualitative linguistic analysis comparing the textual characteristics of the stories generated by the different models. In the first experiment, we asked 68 participants to rate short stories generated by the models and humans along dimensions such as grammaticality, relevance, creativity, and attractiveness. BART Large outperformed human writers in most aspects, except creativity, with an overall score of 2.11 compared to 1.85 for human-written texts—a 14% improvement. In the second experiment, the qualitative analysis revealed that, while GPT-4o exhibited near-perfect internal and external coherence, it tended to produce more predictable narratives, with only 3% of its stories see as novel. In contrast, 15% of BART’s stories were considered novel, indicating a higher degree of creativity despite its smaller model size. This study provides both quantitative and qualitative insights into how model size and fine-tuning influence the balance between creativity, fluency, and coherence in a creative writing tasks.
@@ 1 Introduction
The field of Natural Language Processing (NLP) has seen significant advancements due to Large Language Models (LLMs), which have demonstrated remarkable performance across a variety of tasks, including creative writing Chang et al. (2024); Wang et al. (2024). These models, such as GPT-3.5 and GPT-4 Ouyang et al. (2022); Achiam et al. (2023), can generate human-like text and have set a new standard for many language-related applications. However, these models requires immense computational resources and large datasets, making them resource-intensive. While these large models have excellent performance, recent research has suggested that smaller models, or Small Language Models (SLMs), can achieve competitive results in certain tasks with a much lower computational cost Schick and Schütze (2020); Bilenko (2024); Chen and Varoquaux (2024). This leads to our core research question: RQ0: Can a fine-tuned small language model be competitive in the field of creative, literary writing?
This study explores the capabilities of BART-large Lewis et al. (2019), an SLM, in a creative writing task: generating movie synopses based on a given title. The first goal is to compare the quality of these synopses to those written by humans. To explore this, we conducted a comprehensive study in which we collected over 24,000 manual assessments. Specifically, 68 participants evaluated 60 synopses across five dimensions —readability, understandability, relevance, attractiveness, and creativity— in three different experimental settings. With this dataset we answer the central question: RQ1: How do creative texts produced by SLMs compare with human equivalents in terms of readability, understandability, relevance, attractiveness, and creativity? Our findings show that, surprisingly, BART-large outperformed humans across all quality dimensions except creativity.
@@
Having conducted the experiments with human evaluators comparing texts written by an SLM and texts written by human evaluators, we arrive at our second goal of our research: to study the impact of model size on creative tasks. Thus, we arrive to the last research question: RQ4: Do larger language models directly lead to more creative texts? For this purpose, we conducted a qualitative linguistic analysis that zooms in on the weak and strong points of each model regarding the generation of creative texts. The results of the analysis indicate that, while larger models such as GPT-4o produce more consistent and coherent text, they also tend to follow more predictable, formulaic patterns. In contrast, BART-large, despite being a smaller model, generated more novel and surprising content. Specifically, 15% of the synopses produced by BART-large were deemed creatively novel, compared to only 3% for GPT-4. This suggests that larger models may sacrifice creative flexibility for consistency and fluency, a potential trade-off between model size and creative output.
@@
This research highlights the potential of smaller models for creative tasks, demonstrating that SLMs can sometimes can be competitive with larger models in areas like creativity. This opens up new possibilities for more efficient, task-specific models that balance the strengths of both large and small models.
2 Releated Work
@@
Researchers have explored the potential of small language models (SLMs) in various natural language processing tasks, including creative writing. Schick and Schütze (2021) demonstrated that SLMs could be effective few-shot learners through prompt-based learning and optimization techniques. Eldan and Li (2023) introduced TinyStories, showing that models with fewer than 10 million parameters can generate coherent short stories, challenging the notion that larger models are necessary for fluent text generation. For more comprehensive information on the role of small models in the era of large language models, see the survey by Chen and Varoquaux (2024).
However, evaluating the creativity of language models remains a complex challenge. Chakrabarty et al. (2024) introduced the Torrance Test of Creative Writing to objectively assess creativity, finding that LLM-generated stories pass significantly fewer creativity tests compared to those written by professional authors.
Marco et al. (2024) conducted a study titled "Pron vs. Prompt," where they evaluated a top LLM, GPT-4, against a world-class fiction author, Patricio Pron, with assessments conducted by literary critics. Their findings indicate that LLMs are still far from matching the creative writing skills of top human authors, particularly in terms of originality and literary quality. For a detailed overview of creativity and machine learning, refer to the survey by Franceschelli and Musolesi (2022).
Our work takes the methodology of "Pron vs. Prompt" as a starting point but differs in scope and focus. While the previous study evaluated a top LLM against a top writer with evaluations by literary critics, we aim to evaluate a small language model (BART Large) against average human writers, with assessments conducted by average readers instead of experts. By comparing the performance of BART Large to both human writers and larger models like GPT-4o, we seek to determine whether an SLM can be competitive in creative literary writing among general audiences.
@@ 3 Experimental Design
To answer our research questions, we have designed two different experiments. The first one is an evaluation with humans that tries to quantify how far the texts written by our SLM are from those written by humans. The second experiment consists of qualitatively analyzing the linguistic similarities and differences between the SLM texts and the most popular LLMs right now: GPT-3.5111In July 2024, OpenAI announced that it would replace gpt-3.5 with gpt-4o-mini in the ChatGPT web version. By then this research had already been completed. However, we decided to keep the results of GPT-3.5 because it is still available for access through the API, but, above all, because of its proximity in performance with open source models as Mixtral Jiang et al. (2024) or Gemma Mesnard et al. (2024) that are still widely used. and GPT-4o. The task consists of, given a potential movie title, writing an imaginary synopsis for a movie with that title.
@@
For our main experiment we have used the default BART-large pre-trained model (Lewis et al., 2019) and we have fine-tuned it to the down-stream task of synopsis generation from titles. During generation, we use sampling and beam search with the default parameters of BART HuggingFace model configuration [^2] https://huggingface.co/facebook/bart-large/blob/main/config.json.
@@
We ask the assessor to evaluate a number of quality aspects of the text. We have run three experiments with a similar setup (see "Variants" below). For each experiment, we had two quizzes, each consisting of 60 synopses: half (30) human-written, and half (30) generated by our system. We decided to use this number of synopses because a sample size of 30 is large enough for statistical significance tests, and over 60 synopses would cause fatigue in our assessors. The synopses with human written synopses in quiz A have model synopses in quiz B, and vice versa. Half of the participants were randomly assigned to quiz A and half were randomly assigned to quiz B. The synopses are displayed in random order for each subject.
@@
In order to do this, we have first asked GPT-3.5 and GPT-4o to generate a synopsis for each of the titles in our dataset, simply using the prompt “Invent a synopsis for a movie entitled [MOVIE TITLE]”. Then, we have performed a comparative linguistic study of the synopses generated by the three LLMs. Our results shed some light on the comparative capabilities of small-size, fined tuned model and state-of-the-art models in zero-shot mode.
First, we have manually examined all AI-generated synopses (BART, GPT-3.5 and GPT-4o), and conclude that there are five salient properties that are typical of AI-generated synopses (See Table 3): repetitiveness, recurrent themes, coherence with external facts, internal coherence and surprising associations. Then, each synopsis has been analyzed and annotated by a linguist, taking into account the proposed properties. This provides a qualitative characterization of BART and GPTs synopsis.
@@
Table 1 shows the average scores for each quality aspect of the human vs model-generated synopses, along with the statistical significance of the differences found. The table allows to answer our two first Research Questions.
@@
Does the apprentice beats the master? The fact that a SLM, trained to mimic human-written texts, can outperform humans in such a creative task is notable, as it suggests that they are capable of producing better-than-human outputs even by merely imitating human examples. A possible explanation for the higher average score of the SLM could be that, in the learning process, generalization helps the SLM avoid gross mistakes, and therefore produce more homogeneous text. Then, it would get better averages just by avoiding lowest scores. But this hypothesis is incorrect: although the standard deviation of humans in overall score is higher than in the SLM (0.461 vs 0.355) – which is consistent with the hypothesis – the SLM actually gets more overall 5’s (best possible score) than humans: not only the average is better, but the SLM writes more high quality synopses than the humans. For a more in-depth statistical analysis see Appendix C.
4.2 RQ2: How relevant are the expectations and biases of the reader with respect to AI vs human authors?
@@
In the "revealed author" variant, the bias against AI-generated synopses becomes evident. The overall score difference between humans and AI decreased from 0.257 to 0.107, with AI synopses penalized by 6% when the author was revealed. AI synopses scored lower in readability, understandability, and attractiveness, while human scores remained relatively stable, confirming a negative bias against AI-generated content.
In the "AI-only" variant, where assessors believed all synopses were AI-generated, both human and AI synopses were penalized in readability and understandability. Interestingly, attractiveness ratings increased for both, although this effect is harder to interpret. The overall score difference between humans and AI decreased slightly, but AI synopses still outperformed human ones.
These results suggest that revealing the author creates a bias against AI-generated content, impacting readers’ experience, but not enough to change the overall outcome.
@@ 4.3 RQ3: How do readers assess creativity, and how it correlates with other aspects of the texts?
We deliberately chose not to provide readers with a specific definition of creativity, to avoid introducing our own biases into their evaluations. Our readers were free to interpret the term "creativity”, following Colton and Wiggins (2021): ”Computer Creativity is the philosophy, science and engineering of computational systems which, by taking on particular responsibilities, exhibit behaviors that unbiased observers would deem to be creative.“ Once readers completed the evaluation, we asked them what factors in the synopses influenced their quantitative assessment of creativity. The reason is that we are focused on whether SLM texts give the appearance of creativity to readers in general (something that is measurable), not whether they are truly creative from an artist’s perspective. After completing the experiment, participants were asked what factors influenced their creativity scores.
@@ 5 SLM vs LLMs: Results of the Linguistic Analysis
Up to this point, we have focused on the comparison between human-generated texts and those produced by BART, where BART demonstrated superior performance in most evaluated dimensions. Having established this, we now turn our attention to a more detailed examination of BART in contrast to larger language models (LLMs), such as GPT-3.5 and GPT-4o. This analysis will explore the distinct linguistic features of these models, with a particular emphasis on how the larger LLMs tend to surpass BART in various aspects, especially in the case of GPT-4o. For concrete examples of the linguistic features discussed here, see Appendix E and D.
@@ 5.3 Coherence with External Facts
When focusing on synopses that do mention external facts, both BART and GPT-3.5 show low reliability, with BART being accurate 62.5% of the time regarding places and dates, and GPT-3.5 at 60%. In references to specific historical events, BART is accurate 50% of the time, while GPT-3.5 only reaches 30%. GPT-4o, on the other hand, does not display any inconsistencies in these areas, but only 5% of its synopses reference historical events. These inconsistencies do not heavily impact average manual assessments because the synopses rarely mention historical places and dates.
@@ 6 Conclusions
This study highlights the potential of small language models (SLMs) in creative writing tasks, specifically when generating short stories. Our experiments show that fine-tuned SLMs like BART-large can outperform average humans in many aspects such as readability, understandability, relevance, and informativeness. However, humans still retain a slight advantage in creativity, although the difference is not statistically significant in our data.
Our analysis also reveals that reader biases toward AI authorship negatively influence assessments of AI-generated texts, particularly in creative domains. Despite these biases, SLMs still produced more attractive synopses overall compared to human writers, suggesting that smaller models can offer a competitive edge in creative applications, especially when tasked with generating shorter texts.
Comparing the SLM with larger language models (LLMs) – GPT-3.5 and GPT-4o in our experiments –, our qualitative linguistic analysis revealed that while larger models produce more consistent and coherent texts, they tend to generate more predictable and formulaic narratives. GPT-4o, despite its near-perfect internal and external coherence, produced stories that were seen as novel only 3% of the time. In contrast, the SLM generated novel content in 15% of its stories, indicating a higher degree of creativity despite its smaller size. This suggests a trade-off between model size and creative flexibility, where larger models may prioritize consistency over originality.
These findings emphasize that SLMs can be competitive with both (average) humans and larger models in creative writing tasks, particularly when the task benefits from creative flexibility over strict consistency. It underscores that a Large Language Model is not always necessary, and that fine-tuned smaller models can be more suitable and efficient for specific applications. This opens up possibilities for developing more efficient, task-specific models that balance creativity, fluency, and coherence without the computational overhead of larger models. [^3] All annotated data and model weights are available on GitHub: https://github.com/annon-submission/slm-creativity.
@@
SLM have clearly received better average scores than humans in our experiment. Initially, we did not expect such a strong result. Although language models are trained with more text than a human can read in a lifetime, their generalization abilities are still weak in comparison with humans, and they do not connect language with real world knowledge. Our qualitative analysis, in fact, confirms the limitations of language models. What is, then, the correct interpretation of our results? Do they actually mean that SLM or LLMs can already perform certain creative tasks better than humans? In order to properly answer this question, let us review our experimental setting and how it constrains the scope of our results.
@@
As we have noted earlier, our task requires writing short texts, where internal coherence is less of a challenge. We chose it because it is well known that SLM struggle to keep coherence, and we wanted to measure their creative writing abilities in isolation from that aspect. We cannot, therefore, extrapolate our results to longer texts.
Also, producing movie synopsis is a special kind of writing task. Synopses in the training set have not been created from a movie title, of course, but from the movie itself, and their goal is to inspire interest in the reader without revealing too much about the plot. It is a good choice for evaluation because it is a constrained task, and because there are many examples to learn from. But it is not a canonical creative writing task, and more work is needed before extrapolating our results to other tasks.
The human writers we are comparing with
Although the set of movies used in our experiment with readers were chosen randomly, we had a few constraints. One of them is that the movie must have less than 1,000 reviews on IMDB, which reduces the possibility of our readers being familiarized with the actual movie (in order to avoid and biased evaluations). A side effect of this constraint is that less popular movies may correlate with lesser quality movies. We have checked our dataset with IMDB, and the average rating of the movies in the training set is 6.2 (with a median of 6.4); in the 60 movies chosen for the quiz, on the other hand, the average is 5.7 (median 5.9). The difference is not large, but it may be a partial explanation of why the SLM is able to apparently improve the synopses used in its learning process.
@@
Our assessors are not professional creative writers or critics: we use a popularity criteria in our evaluation design (Do target readers enjoy the synopses?) instead of a professional criteria (How do critics assess the synopsis?). Both types of assessment are legitimate and complementary, and also often contradictory (for instance, blockbusters are usually much more appreciated by the audience than by the critics). In our experiment, we seemed to have reached (and surpassed) the threshold in which machines are able to match human writings in terms of popularity criteria; and this means that popularity criteria alone may not be enough anymore to perform this type of evaluation; and we may need to dive into the (turbulent) waters of how to properly assess the output of a creative process.

### v2

@@ Abstract
In this paper, we evaluate the creative fiction writing abilities of a fine-tuned small language model (SLM), BART-large, and compare its performance to human writers and two large language models (LLMs): GPT-3.5 and GPT-4o. Our evaluation consists of two experiments: (i) a human study in which 68 participants rated short stories from humans and the SLM on grammaticality, relevance, creativity, and attractiveness, and (ii) a qualitative linguistic analysis examining the textual characteristics of stories produced by each model. In the first experiment, BART-large outscored average human writers overall (2.11 vs. 1.85), a 14% relative improvement, though the slight human advantage in creativity was not statistically significant. In the second experiment, qualitative analysis showed that while GPT-4o demonstrated near-perfect coherence and used less cliche phrases, it tended to produce more predictable language, with only 3% of its synopses featuring surprising associations (compared to 15% for BART). These findings highlight how model size and fine-tuning influence the balance between creativity, fluency, and coherence in creative writing tasks, and demonstrate that smaller models can, in certain contexts, rival both humans and larger models. [^1] All annotated data and model weights are available on GitHub: https://github.com/grmarco/slm-creativity.
@@ 1 Introduction
The field of Natural Language Processing (NLP) has seen significant advancements due to Large Language Models (LLMs), which have demonstrated remarkable performance across a variety of tasks, including creative writing Chang et al. (2024); Wang et al. (2024). These models, such as GPT-3.5 and GPT-4 Ouyang et al. (2022); Achiam et al. (2023), can generate human-like text and have set a new standard for many language-related applications. However, they require immense computational resources and large datasets, making them resource-intensive. While these large models have excellent performance, recent research has suggested that smaller models, or Small Language Models (SLMs), can achieve competitive results in certain tasks with a much lower computational cost Schick and Schütze (2020); Bilenko (2024); Chen and Varoquaux (2024). This leads to our core research question: RQ0: Can a fine-tuned small language model be competitive in the field of creative, literary writing?
This study explores the capabilities of BART-large Lewis et al. (2019), an SLM, in a creative writing task: generating movie synopses based on a given title. The first goal is to compare the quality of these synopses to those written by humans. To explore this, we conducted a comprehensive study in which we collected over 24,000 manual assessments. Specifically, 68 participants evaluated 60 synopses across five dimensions —readability, understandability, relevance, attractiveness, and creativity— in three different experimental settings. With this dataset, we answer the central question: RQ1: How do creative texts produced by SLMs compare with human equivalents in terms of readability, understandability, relevance, attractiveness, and creativity? Our findings show that, surprisingly, BART-large outperformed humans across all quality dimensions except creativity.
@@
Having conducted the experiments with human evaluators comparing texts written by an SLM and texts written by human authors, we arrive at our second goal of our research: to study the impact of model size on creative tasks. Thus, we arrive at the last research question: RQ4: Do larger language models directly lead to more creative texts? For this purpose, we conducted a qualitative linguistic analysis that zooms in on the weak and strong points of each model regarding the generation of creative texts. The results of the analysis indicate that, while larger models such as GPT-4o produce more consistent and coherent text, they also tend to follow more predictable, formulaic patterns. In contrast, BART-large, despite being a smaller model, generated more novel and surprising content. Specifically, 15% of the synopses produced by BART-large were deemed creatively novel, compared to only 3% for GPT-4. This suggests that larger models reproduce natural language patterns more accurately, achieving greater consistency and fluency but often generating more predictable and formulaic outputs. In contrast, smaller models like BART-large, since they do not replicate these patterns as precisely, are more likely to deviate from predictable structures, leading to greater flexibility and originality. This highlights how the inherent limitations of smaller models can sometimes foster more creative and unexpected outputs.
@@
This research highlights the potential of smaller models for creative tasks, demonstrating that SLMs can sometimes be competitive with larger models in areas like creativity. This opens up new possibilities for more efficient, task-specific models that balance the strengths of both large and small models.
2 Related Work
@@
Early research has shown that small language models (SLMs) can perform well in creative text generation. For example, Ravi et al. (2024) found that smaller models can produce humor when given proper guidance. Similarly, Schick and Schütze (2021) demonstrated that SLMs work effectively as few-shot learners with prompt-based techniques. Eldan and Li (2023) introduced TinyStories, showing that models with fewer than 10 million parameters can still generate coherent short narratives, challenging the belief that larger models are always needed for fluent and coherent output. For a more detailed overview of small models, see Chen and Varoquaux (2024).
Work on how model size and perceived authorship influence reader evaluations has produced mixed results. In poetry, Porter and Machery (2024) found that a large language model can sometimes be preferred over human-authored poems, especially if readers are unaware of the poem’s origin. In contrast, Marco et al. (2024), who compared a leading LLM (GPT-4) with a renowned human author and used expert literary critics as evaluators, concluded that the top human author was clearly favored, particularly for originality and literary quality. These findings suggest that preferences depend on the evaluator’s expertise, the model’s scale, and the author’s standing.
Our study explores a different combination: we compare a small language model (BART-large) against average human writers, evaluated by general readers. By varying the information provided about the text’s origin—explicitly stating that it is AI-generated, stating that it is human-written, or withholding the information—we investigate whether a smaller model can match or surpass the performance of both larger models and human authors under these conditions.
Assessing creativity itself remains a complex challenge. Tests like the Torrance Test of Creative Writing Chakrabarty et al. (2024) show that LLM-generated texts often do not meet the creative standards set by professional authors. For more information on creativity in machine learning, see Franceschelli and Musolesi (2022). In line with these findings, our study considers how small models fit into this broader landscape, examining how model size, author expertise, and audience perception interact.
@@ 3 Experimental Design
To answer our research questions, we have designed two different experiments. The first one is an evaluation with humans that tries to quantify how far the texts written by our SLM are from those written by humans. The second experiment consists of qualitatively analyzing the linguistic similarities and differences between the SLM texts and the most popular LLMs right now: GPT-3.5222In July 2024, OpenAI announced that it would replace gpt-3.5 with gpt-4o-mini in the ChatGPT web version. By then this research had already been completed. However, we decided to keep the results of GPT-3.5 because it is still available for access through the API, but, above all, because of its proximity in performance with open source models as Mixtral Jiang et al. (2024) or Gemma Mesnard et al. (2024) that are still widely used. and GPT-4o. The task consists of, given a potential movie title, writing an imaginary synopsis for a movie with that title.
@@
For our main experiment we used the default BART-large pre-trained model (Lewis et al., 2019). Recent work in humor generation Ravi et al. (2024) shows that BART remains a strong model for creative writing tasks. We fine-tuned it for synopsis generation from titles. We rely on sampling and beam search with the default BART HuggingFace configuration [^3] https://huggingface.co/facebook/bart-large/blob/main/config.json, without adjusting the temperature parameter. Although it is often suggested that increasing temperature could foster greater creative variability, recent research Peeperkorn et al. (2024) indicates that higher temperatures mainly introduce randomness rather than yielding truly more creative outputs. Consequently, we maintain a fixed temperature setting. This ensures that observed differences in textual quality or reader perception stem from the model’s inherent capabilities and our experimental conditions rather than from artificial variability.
@@
We ask the assessor to evaluate a number of quality aspects of the text. We have run three experiments with a similar setup (see "Variants" below). For each experiment, we had two quizzes, each consisting of 60 synopses: half (30) human-written, and half (30) generated by our system. We decided to use this number of synopses because a sample size of 30 is large enough for statistical significance tests, and over 60 synopses would cause fatigue in our assessors. The synopses with human-written synopses in quiz A have model synopses in quiz B, and vice versa. Half of the participants were randomly assigned to quiz A and half were randomly assigned to quiz B. The synopses are displayed in random order for each subject.
@@
In order to do this, we have first asked GPT-3.5 and GPT-4o to generate a synopsis for each of the titles in our dataset, simply using the prompt “Invent a synopsis for a movie entitled [MOVIE TITLE]”. Then, we performed a comparative linguistic study of the synopses generated by the three LLMs. Our results shed some light on the comparative capabilities of small-size, fined tuned model and state-of-the-art models in zero-shot mode.
First, we have manually examined all AI-generated synopses (BART, GPT-3.5 and GPT-4o), and conclude that there are five salient properties that are typical of AI-generated synopses (See Table 3): repetitiveness, recurrent themes, coherence with external facts, internal coherence and surprising associations. Then, each synopsis has been analyzed and annotated by a linguist, taking into account the proposed properties. This provides a qualitative characterization of BART and GPTs synopses.
@@
Table 1 shows the average scores for each quality aspect of the human vs. model-generated synopses, along with the statistical significance of the differences found. The table allows us to answer our first two Research Questions.
@@
Does the apprentice beat the master? The fact that a SLM, trained to mimic human-written texts, can outperform humans in such a creative task is notable, as it suggests that they are capable of producing better-than-human outputs even by merely imitating human examples. A possible explanation for the higher average score of the SLM could be that, in the learning process, generalization helps the SLM avoid gross mistakes, and therefore produce more homogeneous text. Then, it would get better averages just by avoiding lowest scores. However, this hypothesis is incorrect: although the standard deviation of humans in overall score is higher than in the SLM (0.461 vs. 0.355) – which is consistent with the hypothesis – the SLM actually gets more overall 5’s (best possible score) than humans: not only the average is better, but the SLM writes more high quality synopses than the humans. For a more in-depth statistical analysis see Appendix C.
4.2 RQ2: How relevant are the expectations and biases of the reader with respect to AI vs. human authors?
@@
In the "revealed author" variant, the bias against AI-generated synopses becomes evident. The overall score difference between humans and AI decreased from 0.257 to 0.107, with AI synopses penalized by 6% when the author was revealed. AI synopses scored lower in readability, understandability, and attractiveness, while human scores remained relatively stable. This suggests a clear negative bias against AI-generated content when directly compared to human authors, as the explicit framing highlights AI as a competitor.
In contrast, in the "AI-only" variant, where assessors believed all synopses were AI-generated, the dynamic changes. Both human and AI synopses were penalized in readability and understandability, likely due to preconceived notions about AI’s limitations. However, attractiveness ratings increased for both groups, and some assessors expressed surprise or admiration for the perceived quality of the synopses. This "wow" effect may stem from the novelty or unexpected competence of AI in generating coherent and engaging content, especially when readers are not directly comparing it to human output. Interestingly, this effect suggests that AI’s perceived performance can benefit when it is not framed as a direct competitor to human creativity.
These findings indicate that revealing the author introduces a bias against AI-generated content, lowering reader evaluations, particularly in direct comparisons with human authors. However, when AI is presented as the sole source of content, some biases may transform into a positive re-evaluation of its capabilities, highlighting the nuanced interplay between reader expectations and the framing of authorship.
@@ 4.3 RQ3: How do readers assess creativity, and how it correlates with other aspects of the texts?
We deliberately chose not to provide readers with a specific definition of creativity, to avoid introducing our own biases into their evaluations. Our readers were free to interpret the term "creativity”, following Colton and Wiggins (2021): “Computer Creativity is the philosophy, science and engineering of computational systems which, by taking on particular responsibilities, exhibit behaviors that unbiased observers would deem to be creative.“ Once readers completed the evaluation, we asked them what factors in the synopses influenced their quantitative assessment of creativity. The reason is that we are focused on whether SLM texts give the appearance of creativity to readers in general (something that is measurable), not whether they are truly creative from an artist’s perspective. After completing the experiment, participants were asked what factors influenced their creativity scores.
@@ 5 SLM vs LLMs: Results of the Linguistic Analysis
Up to this point, we have focused on the comparison between human-generated texts and those produced by BART, where BART demonstrated superior performance in most evaluated dimensions. Having established this, we now turn our attention to a more detailed examination of BART in contrast to larger language models (LLMs), such as GPT-3.5 and GPT-4o. This analysis will explore the distinct linguistic features of these models, with a particular emphasis on how the larger LLMs tend to surpass BART in various aspects, especially in the case of GPT-4o. For concrete examples of the linguistic features discussed here, see Appendices E and D.
@@ 5.3 Coherence with External Facts
When focusing on synopses that do mention external facts, both BART and GPT-3.5 show low reliability, with BART being accurate 62.5% of the time regarding places and dates, and GPT-3.5 at 60%. In references to specific historical events, BART is accurate 50% of the time, while GPT-3.5 only reaches 30%. GPT-4o, on the other hand, does not display any inconsistencies in these areas, but only 5% of its synopses refer historical events. These inconsistencies do not heavily impact average manual assessments because the synopses rarely mention historical places and dates.
@@ 6 Conclusions
This study highlights the potential of small language models (SLMs) in creative writing tasks, particularly in generating short stories. Our experiments indicate that fine-tuned SLMs like BART-large are rated more favorably than average human writers by general readers in several aspects, including readability, understandability, relevance, and informativeness. However, when it comes to creativity, humans maintain a slight perceived edge, though the difference observed in our data is not statistically significant.
Interestingly, our findings suggest that it does not take a highly sophisticated large language model (LLM) to meet or even surpass human performance in many dimensions of creative writing, particularly when evaluated by general audiences. SLMs, despite their smaller size and lower computational requirements, demonstrated a strong ability to generate texts that resonated with readers, challenging assumptions about the necessity of larger models in this domain.
Our results also highlight the significant role of biases in shaping reader perceptions of AI-generated text. When AI is explicitly identified as the author, reader assessments tend to penalize its perceived creativity and overall quality. However, in scenarios where the AI is not competing directly against human authors or is evaluated anonymously, its output often receives more favorable ratings. This effect underlines how context and framing influence the reception of AI-generated content.
Comparing the SLM with larger language models (LLMs) —GPT-3.5 and GPT-4o in our experiments— our qualitative linguistic analysis revealed that while larger models produce more consistent and coherent texts, they often generate more predictable and formulaic narratives. GPT-4o, despite its near-perfect internal and external coherence, produced stories that were seen as novel only 3% of the time. In contrast, the SLM generated novel content in 15% of its stories, indicating a higher degree of creativity despite its smaller size. This suggests a trade-off between model size and creative flexibility, where larger models may prioritize consistency over originality.
These findings emphasize that SLMs can be competitive with both (average) humans and larger models in creative writing tasks, particularly when the task benefits from creative flexibility over strict consistency. It underscores that a Large Language Model is not always necessary, and that fine-tuned smaller models can be more suitable and efficient for specific applications. This opens up possibilities for developing more efficient, task-specific models that balance creativity, fluency, and coherence without the computational overhead of larger models.
@@
SLMs have clearly received better average scores than humans in our experiment. Initially, we did not expect such a strong result. Although language models are trained with more text than a human can read in a lifetime, their generalization abilities are still weak in comparison with humans, and they do not connect language with real world knowledge. Our qualitative analysis, in fact, confirms the limitations of language models. What, then, is the correct interpretation of our results? Do they actually mean that SLM or LLMs can already perform certain creative tasks better than humans? In order to properly answer this question, let us review our experimental setting and how it constrains the scope of our results.
@@
As we have noted earlier, our task requires writing short texts, where internal coherence is less of a challenge. We chose it because it is well known that SLMs struggle to maintain coherence, and we wanted to measure their creative writing abilities without that factor influencing the results. We cannot, therefore, extrapolate our results to longer texts.
Also, producing a movie synopsis is a special kind of writing task. The synopses in the training set were not created from just a movie title, of course, but from the movie itself. Their goal is to spark the reader’s interest without revealing too much of the plot. It is a good choice for evaluation because it is a constrained task, and because there are many examples to learn from. But it is not a canonical creative writing task, and more work is needed before extrapolating our results to other tasks.
The human writers we are comparing against
Although the set of movies used in our experiment with readers were chosen randomly, we had a few constraints. One of them is that the movie must have less than 1,000 reviews on IMDb, which reduces the possibility of our readers being familiarized with the actual movie (in order to avoid any biased evaluations). A side effect of this constraint is that less popular movies may correlate with lesser quality movies. We have checked our dataset with IMDb, and the average rating of the movies in the training set is 6.2 (with a median of 6.4); in the 60 movies chosen for the quiz, on the other hand, the average is 5.7 (median 5.9). The difference is not large, but it may be a partial explanation of why the SLM is able to apparently improve the synopses used in its learning process.
@@
Our assessors are not professional creative writers or critics; instead, we use a popularity criterion in our evaluation design (Do target readers enjoy the synopses?) instead of a professional criteria (How do critics assess the synopsis?). Both types of assessment are legitimate and complementary, and also often contradictory (for instance, blockbusters are usually much more appreciated by the audience than by the critics). In our experiment, we seemed to have reached (and surpassed) the threshold in which machines are able to match human writings in terms of popularity criteria; and this suggests that relying solely on popularity may no longer suffice for this type of evaluation. We may need to venture into the (turbulent) waters of how to properly assess the outputs of a creative process.
@@
Acknowledgements
This work has been financed by the European Union (NextGenerationEU funds) through the “Plan de Recuperación, Transformación y Resiliencia”, by the Ministry of Economic Affairs and Digital Transformation and by UNED University. However, the points of view and opinions expressed in this document are solely those of the authors and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be considered responsible for them.
Guillermo Marco’s work was funded by Spanish government Ph.D. research grant (Ministerio de Universidades) FPU20/07321 and a scholarship of the Madrid City Council for the Residencia de Estudiantes (Course 2023-2024).